# FUSION: Forecast-Embedded Agent Scheduling with Service Incentive Optimization over Distributed Air-Ground Edge Networks

### v2

@@ I Introduction
II Related Work
II-A Service Provision in UAV-Assisted Edge Networks
@@
Although the above studies offer valuable insights, they predominantly presume that UAVs are always nearby and can provide immediate service, often without explicitly modeling their underlying deployment units or origins. As a result, the joint treatment of the four challenges summarized in Sec. I remains largely unexplored, an effort undertaken in this work.
II-B Auction-Driven and Game-Theoretic Service Provisioning
III Overview and Modeling
III-A Overview of Our Methodology
@@
Under this two-sided demand-supply structure, SDs compete for reliable computing services, while SPs collaboratively offer computing resources under their capacity and operational limitations. Nevertheless, in such a dynamic and distributed ecosystem, ESs face uncertain service demand: the temporal and spatial distribution of service request peaks across ESs is highly stochastic, and the composition ratio between HUs and MUs in the service requests also varies over time. To capture the time-varying behavior, a practical transaction is discretized into $T$ time-slots, with index $t$, i.e., $t\in\left\{0,1,2,...,T\right\}$. Further, to enable resource provisioning between SDs and SPs, we adopt a two-stage resource allocation framework that explicitly separates an offline stage from an online stage. From the ES perspective, the offline stage operates before the actual trading round, when only historical records and statistical demand patterns are available. At this stage (detailed in Sec. IV), the identities and tasks of HUs and MUs in the next trading round are unknown, making it impossible to pre-plan precise AP-ES matchings. Accordingly, in the offline stage of Fig. 1, we train Pro-LNN to forecast future service demand for each ES, and, based on these predictions, jointly design AP-to-ES visiting routes, UAV deployment plans, and auction-based contracts between ESs and APs (i.e., Off-AIC2).
Figure 1: Schematic of FUSION over an air–ground integrated network.
In contrast, the online stage starts (detailed in Sec. V) when real computing tasks arrive and the actual identities and locations of SDs become known. With full knowledge of active participants and their instantaneous task requirements, the system performs real-time task–resource scheduling. In this stage, we model the interactions among heterogeneous SPs and SDs as a potential game and construct an ordinal potential that captures the social welfare. Each SD then updates its task offloading decision and each SP adjusts its service admission in a fully distributed manner, where the resulting potential-guided best-response dynamics converge to a pure-strategy NE, yielding an online and congestion-aware provisioning of resources.
III-B Modeling of SDs, SPs, and Contracts
@@ III-B2 Modeling of SPs
III-B3 Modeling of Offline Contracts
@@
IV Offline Stage: Forecasting-Driven Contracting and AP Deployment
@@
IV-A Utilities of ESs, AP, and Auctioneer
IV-A1 Utility of ESs
@@ IV-A3 Utility of the Auctioneer
IV-B Problem of Interest in Offline Stage of FUSION
@@
IV-C Forecasting, Agent Deployment, and Contracting
We next present the offline stage that couples demand forecasting, AP deployment, and long-term contracting, with detailed interaction flow moved to Appx. B-A. In what follows, we describe the three modules, i.e., Pro-LNN-based demand forecasting, eACO-VRP-based AP deployment, and the Off-AIC2 contracting mechanism.
IV-C1 Design of Proactive Spatio-Temporal Demand Prediction via Liquid Neural Networks (Pro-LNN)
@@
Let $N_{i}^{\mathrm{Dem},(\nu)}$ denote the task demand of ES $s_{i}$ at the $\nu$-th trading round. We form an input vector $\mathbf{z}_{i}^{(\nu)}=\big[N_{i}^{\mathrm{Dem},(\nu)},\mathbf{x}_{i}^{(\nu)}\big]$, where $\mathbf{x}_{i}^{(\nu)}$ stacks optional exogenous features at round $\nu$, such as time-of-day and day-of-week indicators, ES-type and capacity descriptors, and local mobility/load statistics extracted from recent logs. Given a length-$L$ history of past observations $\{\mathbf{z}_{i}^{(\nu-L+1)},\ldots,\mathbf{z}_{i}^{(\nu)}\}$, Pro-LNN encodes this sequence via a liquid cell with neuron-wise time constants, integrates the measurements into a hidden state $\mathbf{h}_{i}^{(\nu)}$, and then applies a shallow readout network to produce an $H$-step forecast $\{\hat{N}_{i}^{\mathrm{Dem},(\nu+1)},\ldots,\hat{N}_{i}^{\mathrm{Dem},(\nu+H)}\}$. In a nutshell, Pro-LNN is trained with a composite loss, which is an $H$-step MSE between the forecasts and the ground-truth demands, and we additionally use a multi-step consistency regularizer that rolls the predictor forward in closed-loop over $H$ steps and penalizes the accumulated prediction errors at future trading rounds. These forecasts are computed offline before each trading round and then fed into the eACO-VRP module for ES recommendation and AP route pre-planning, as well as the auction module for ES–AP contract design. The detailed update equations and training pseudocode of our Pro-LNN are deferred to Appx. B-B to avoid detracting from the main discussion.
IV-C2 Enhanced Ant Colony Optimization for Vehicle Route Planning (eACO-VRP)

### v3

@@ 1 Introduction
2 Related Work
2.1 Service Provision in UAV-Assisted Edge Networks
@@
Although the above studies offer valuable insights, they predominantly presume that UAVs are always nearby and can provide immediate service, often without explicitly modeling their underlying deployment units or origins. As a result, the joint treatment of the four challenges summarized in Sec. 1 remains largely unexplored, an effort undertaken in this work.
@@
2.2 Auction-Driven and Game-Theoretic Service Provisioning
3 Overview and Modeling
Figure 1: Schematic of FUSION over an air–ground integrated network.
3.1 Overview of Our Methodology
@@
Under this two-sided demand-supply structure, SDs compete for reliable computing services, while SPs collaboratively offer computing resources under their capacity and operational limitations. Nevertheless, in such a dynamic and distributed ecosystem, ESs face uncertain service demand: the temporal and spatial distribution of service request peaks across ESs is highly stochastic, and the composition ratio between HUs and MUs in the service requests also varies over time. To capture the time-varying behavior, a practical transaction is discretized into $T$ time-slots, with index $t$, i.e., $t\in\left\{0,1,2,...,T\right\}$. Further, to enable resource provisioning between SDs and SPs, we adopt a two-stage resource allocation framework that explicitly separates an offline stage from an online stage. From the ES perspective, the offline stage operates before the actual trading round, when only historical records and statistical demand patterns are available. At this stage (detailed in Sec. 4), the identities and tasks of HUs and MUs in the next trading round are unknown, making it impossible to pre-plan precise AP-ES matchings. Accordingly, in the offline stage of Fig. 1, we train Pro-LNN to forecast future service demand for each ES, and, based on these predictions, jointly design AP-to-ES visiting routes, UAV deployment plans, and auction-based contracts between ESs and APs (i.e., Off-AIC2).
In contrast, the online stage starts (detailed in Sec. 5) when real computing tasks arrive and the actual identities and locations of SDs become known. With full knowledge of active participants and their instantaneous task requirements, the system performs real-time task–resource scheduling. In this stage, we model the interactions among heterogeneous SPs and SDs as a potential game and construct an ordinal potential that captures the social welfare. Each SD then updates its task offloading decision and each SP adjusts its service admission in a fully distributed manner, where the resulting potential-guided best-response dynamics converge to a pure-strategy NE, yielding an online and congestion-aware provisioning of resources.
3.2 Modeling of SDs, SPs, and Contracts
@@ 3.2.2 Modeling of SPs
3.2.3 Modeling of Offline Contracts
@@
4 Offline Stage: Forecasting-Driven Contracting and AP Deployment
@@
4.1 Utilities of ESs, AP, and Auctioneer
4.1.1 Utility of ESs
@@ 4.1.3 Utility of the Auctioneer
4.2 Problem of Interest in Offline Stage of FUSION
@@
4.3 Forecasting, Agent Deployment, and Contracting
We next present the offline stage that couples demand forecasting, AP deployment, and long-term contracting. The overall interaction flow of the proposed Pro-LNN + eACO-VRP + Off-AIC2 framework as follow:
In what follows, we describe the three modules, i.e., Pro-LNN-based demand forecasting, eACO-VRP-based AP deployment, and the Off-AIC2 contracting mechanism.
4.3.1 Design of Proactive Spatio-Temporal Demand Prediction via Liquid Neural Networks (Pro-LNN)
@@
Let $N_{i}^{\mathrm{Dem},(\nu)}$ denote the task demand of ES $s_{i}$ at the $\nu$-th trading round. We form an input vector $\mathbf{z}_{i}^{(\nu)}=\big[N_{i}^{\mathrm{Dem},(\nu)},\mathbf{x}_{i}^{(\nu)}\big]$, where $\mathbf{x}_{i}^{(\nu)}$ stacks optional exogenous features at round $\nu$, such as time-of-day and day-of-week indicators, ES-type and capacity descriptors, and local mobility/load statistics extracted from recent logs. Given a length-$L$ history of past observations $\{\mathbf{z}_{i}^{(\nu-L+1)},\ldots,\mathbf{z}_{i}^{(\nu)}\}$, Pro-LNN encodes this sequence via a liquid cell with neuron-wise time constants, integrates the measurements into a hidden state $\mathbf{h}_{i}^{(\nu)}$, and then applies a shallow readout network to produce an $H$-step forecast $\{\hat{N}_{i}^{\mathrm{Dem},(\nu+1)},\ldots,\hat{N}_{i}^{\mathrm{Dem},(\nu+H)}\}$. In a nutshell, Pro-LNN is trained with a composite loss, which is an $H$-step MSE between the forecasts and the ground-truth demands, and we additionally use a multi-step consistency regularizer that rolls the predictor forward in closed-loop over $H$ steps and penalizes the accumulated prediction errors at future trading rounds. These forecasts are computed offline before each trading round and then fed into the eACO-VRP module for ES recommendation and AP route pre-planning, as well as the auction module for ES–AP contract design. The detailed update equations and training pseudocode of our Pro-LNN are deferred to Appx. B.1 to avoid detracting from the main discussion.
4.3.2 Enhanced Ant Colony Optimization for Vehicle Route Planning (eACO-VRP)